A statellite has the mass of 2000kg realease on earth

It revolves 20km above earth find the force?(R=9559km)


1
Expert's answer
2021-04-05T11:07:42-0400

Let's first find the acceleration of the satellite:


"a=\dfrac{GM_E}{(R_E+h)^2},""a=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot5.98\cdot10^{24}\ kg}{(6.37\cdot10^6\ m+20\cdot10^3\ m)^2}=9.77\ \dfrac{m}{s^2}."

Then, we can find the force:


"F=ma=2000\ kg\cdot9.77\ \dfrac{m}{s^2}=19540\ N."